OpenStudy (anonymous):

What is 3x+y=-2 graphed?

OpenStudy (anonymous):

can you put it standard form?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

-3%2Ax-1%2Ay=2Start with the given equation -1%2Ay=2%2B3%2Ax Add 3%2Ax to both sides y=%28-1%29%282%2B3%2Ax%29 Multiply both sides by -1 y=%28-1%29%282%29%2B%281%29%28-3%29x%29 Distribute -1 y=-2-%283%29x Multiply y=-3%2Ax-2 Rearrange the terms y=-3%2Ax-2 Reduce any fractions So the equation is now in slope-intercept form (y=mx%2Bb) where m=-3 (the slope) and b=-2 (the y-intercept) So to graph this equation lets plug in some points Plug in x=-3 y=-3%2A%28-3%29-2 y=9-2 Multiply y=7 Add So here's one point (-3,7)

OpenStudy (anonymous):

Now lets find another point Plug in x=-2 y=-3%2A%28-2%29-2 y=6-2 Multiply y=4 Add So here's another point (-2,4). Add this to our grap

OpenStudy (anonymous):

So from the graph we can see that the slope is -3%2F1 (which tells us that in order to go from point to point we have to start at one point and go down -3 units and to the right 1 units to get to the next point), the y-intercept is (0,-2)and the x-intercept is (-0.666666666666667,0) ,or (-2%2F3,0) . So all of this information verifies our graph. We could graph this equation another way. Since b=-2 this tells us that the y-intercept (the point where the graph intersects with the y-axis) is (0,-2). So we have one point (0,-2)

OpenStudy (anonymous):

So this is the graph of y=-3%2Ax-2 through the points (0,-2) and (1,-5)

OpenStudy (alekos):

or you could just use Desmos or maybe even Google Graph?